# Anatomy of a Digital Bubble: Lessons Learned from the NFT and Metaverse Frenzy

### v1

@@ Abstract
In the past few years, “metaverse” and “non-fungible tokens (NFT)” have become buzzwords, and the prices of related assets have shown speculative bubble-like behavior. In this paper, we attempt to better understand the underlying economic dynamics. To do so, we look at Decentraland, a virtual world platform where land parcels are sold as NFT collections. We find that initially, land prices followed traditional real estate pricing models – in particular, value decreased with distance from the most desirable areas – suggesting Decentraland behaved much like a virtual city. However, these real estate pricing models stopped applying when both the metaverse and NFTs gained increased popular attention and enthusiasm in 2021, suggesting a new driving force for the underlying asset prices. At that time, following a substantial rise in NFT market values, short-term holders of multiple parcels began to take major selling positions in the Decentraland market, which hints that, rather than building a metaverse community, early Decentraland investors preferred to cash out when land valuations became overly inflated. Our analysis also shows that while the majority of buyers are new entrants to the market (many of whom joined during the bubble), liquidity (i.e., parcels) was mostly provided by early adopters selling, which caused stark differences in monetary gains. Early adopters made money – more than 10,000 USD on average per parcel sold – but users who joined later typically made no profit or even incurred losses in the order of 1,000 USD per parcel. Unlike established markets such as financial and real estate markets, newly emergent digital marketplaces are mostly self-regulated. As a result, the significant financial risks we identify indicate a strong need for establishing appropriate standards of business conduct and improving user awareness.
@@ 1 Introduction
The Metaverse is a term originally coined in science fiction in 1992 [36], to denote an extension of the Internet into a single virtual world. With advances in computer graphics and virtual reality hardware, we have seen increasingly realistic attempts to build a metaverse – or at least a number of virtual worlds – over the past couple of decades, with platforms such as Second Life becoming quite popular. Tellingly, in a bid that virtual worlds are about to harness enormous economic potential, in October 2021, Facebook rebranded itself as “Meta,” hoping to become one of the main building blocks of this metaverse [45].
@@
Until recently, most virtual worlds (e.g., Second Life, World of Warcraft, etc.) were centralized, and operated by a single company. Although second-hand markets exist, in traditional, centralized virtual worlds, users willing to purchase wearables, e.g., to decorate their virtual-world avatars, or acquire virtual land, usually have to acquire them from the hosting company.
@@
On the other hand, newer projects, such as Decentraland [^1] https://decentraland.org/. or The Sandbox, [^2] https://www.sandbox.game/en/. have leveraged recent advances in decentralization technologies – e.g., cryptocurrencies that instantiate bankless peer-to-peer payment systems – as a data layer on which vital records of virtual worlds, governed by decentralized anonymous organizations (DAOs), are hosted. Roughly speaking, DAOs attempt to implement a democratic process where governance tokens can be purchased by the general public. Similar to holding company shares, votes are typically weighted by the amount of governance tokens one holds [37]. Likewise, decentralized virtual world items (e.g., wearables, land, and so forth) are often instantiated as “non-fungible tokens” (NFTs), which are also supported by (distributed) blockchain infrastructure rather than a hosting company. Importantly, both governance tokens and NFTs can be traded outside the virtual world platform, e.g., on cryptocurrency exchanges, or third-party NFT marketplaces such as OpenSea. [^3] https://opensea.io.
@@
Decentralized virtual worlds appear to have capitalized on the surge in popularity of both the metaverse and NFTs. Concretely, in late 2021, the renewed interest in the metaverse following Facebook’s announcement, combined with the simultaneous sudden and substantial rise in cryptocurrency prices, caused a massive increase in Decentraland virtual land (parcel) prices, and in the valuation of its governance token called MANA. The median sales price of land in Decentraland rose from less than 1,000 USD in 2020 to roughly 15,000 USD in Q4 of 2021—only to plummet back to 5,000 USD in the subsequent half year. At the time of writing, prices have gone back down to approximately 1,000 USD.
@@
Such volatility has all the hallmarks of an economic bubble. How did the situation affect the wealth of all investors participating in Decentraland—both buyers and sellers? Were certain groups of users disproportionately and/or adversely affected? If so, could regulation help “level the playing field?” Our main contribution is addressing these questions based on a quantitative analysis of online conversations (on Reddit) and of Ethereum blockchain records about Decentraland.
@@
Second, we examine whether traditional (physical) real estate pricing models apply to Decentraland land prices to understand if at some time these prices reflected any traditional notions of the fundamental value of real estate. Decentraland uses a finite Euclidean topology to subdivide the entire world into parcels, each of which is paired to a unique NFT. Similar to holding a property title in the physical world, the owner of that NFT owns the parcel, and, accordingly, can modify its contents (e.g., build something), start a business, or sell it. Different from the physical world, though, Decentraland allows users to instantly teleport anywhere within the virtual world. The question then becomes whether parcel prices follow traditional real-estate price patterns – with prices decreasing as a function of distance from a desirable location – indicating a fundamental feature present in a virtual city.
We show that in the early period when prices were low, parcel prices do in fact appear to follow this class of traditional real-estate models quite closely. Instead, during the period of the rapid rise in land prices, location no longer appears to matter much–investors seem to purchase plots like they purchase other NFTs (e.g., art or collectibles), presumably for speculative purposes. In other words, land prices begin to deviate from the traditional real-estate models, suggesting a bubble-like phenomenon. For this reason, we describe this rise in prices as the “NFT and metaverse bubble.”
Our analysis of market demographics further shows that two distinct groups played major roles during the period we study. New entrants (i.e., those who started trading during the bubble), most of whom appear to be individual investors, were dominantly net-buyer groups (i.e., they bought more than they sold). On the other hand, early adopters provided liquidity (i.e., parcels) to the market and were mostly selling. This caused stark differences in monetary profits: Early adopters reaped 15,000 USD worth of profit per parcel on average. On the other hand, new entrants could rarely profit from the sales – in general, they lost in the order of 1,000 USD per parcel. In short, the bubble led to a disproportionate flow of money from new participants to early adopters.
Digital item marketplaces fostered by the metaverse and NFT are currently subject to lax regulations or left as a grey zone [25, 40, 56]. However, our results highlight significant financial risk to new market entrants, who may have been lured by the promise of riches, when the situation was markedly unstable. These new market entrants may not have had enough information to make rational or well-informed decisions. This result echoes the negative impact of the frenzy surrounding blockchain technology [17] and emphasizes the need for establishing appropriate standards of business conduct and improving user awareness, for example, through education or mandatory risk disclosure.
@@ 2 Related work and background
Public blockchains were originally introduced to realize a secure and transparent ledger system without relying on trusted third parties [49]. However, this technology is now supporting a large volume of mostly speculative cryptocurrency trading [11, 58, 39]. Besides cryptocurrencies, blockchains are also used to support non-fungible tokens (NFTs) [48]. Smart contracts, scripts to automate transactions, make secure and transparent tracking of each token’s ownership possible, by following established NFT standards [24]. In particular, even though all transactions are executed on publicly available ledger systems, NFT ownership is resistant to tampering, without relying on trusted third-parties. Hence, NFTs have been prominently touted as a technology that facilitates the trade of creative works in a secure and transparent manner. Another desirable feature of NFTs is that, being supported by decentralized blockchains, they are platform-independent. In particular, they are not tied to specific seller marketplaces: anybody can buy and sell them. As a result, metaverse platforms that use NFTs to represent in-world items allow these items to be traded outside the platform.
Our hope for this work is to complement the literature that has evidenced the possible negative impact of hype cycles surrounding blockchain technology [17, 37], by providing a quantitative analysis of the economic dynamics of the NFT market for metaverse-related items. More broadly, we hope to contribute to the pressing agenda of better understanding emerging “Web3” businesses [12].
@@ 3 Dataset
This section briefly describes Decentraland first and explains our dataset. Our measurements and analysis is based on three data sources: (1) Ethereum blockchain records, (2) Decentraland public APIs, and (3) Reddit posts.
@@ 3.1 Decentraland dataset
Decentraland is an online platform that has been developed since 2015. Public interest in Decentraland grew in 2017 when developers held an initial coin offering (ICO) of their MANA governance token along with the first land auction. In the auction, participants had the option to bid on parcels of land or stake their MANA towards community-initiated districts. [^5] See https://medium.com/decentraland/the-terraform-event-is-around-the-corner-bdda6a2d4367 and https://github.com/decentraland/districts. Following a second land auction in 2018 and the introduction of developer tools, the official Decentraland client was launched in February 2020. At the time of writing Decentraland has published roughly 93,000 parcels of land as an NFT collection called LAND.
@@
Ether, the utility token for the Ethereum blockchain, has the second-largest market capitalization (after Bitcoin) at the time of writing. [^6] https://coinmarketcap.com/ Ethereum’s main feature is the ability to deploy smart contracts. In a nutshell, smart contracts are programs independently computed by all participants in the (Ethereum) network, who then collectively reach a consensus on the program output. This allows computations to be extremely robust to attacks. Decentraland employs smart contracts as a critical infrastructure. Because all Ethereum transactions, including all interactions with smart contracts, are recorded in a public ledger (the Ethereum blockchain), we can readily explore details of any Decentraland transaction, such as the contract functions that were called, the parties to the transaction, and any fees they might have incurred.
@@
Decentraland provides public APIs to collect parcel data, including parcel property information and ownership transfer history, sorted chronologically. [^10] https://github.com/decentraland/LAND-permissions-graph; https://github.com/decentraland/atlas-server. Every LAND token, including those bundled as EST tokens, has one or more property tags attached to it to denote the type of property the parcel refers to: 1) personally-owned lands can be bought and sold by users; 2) roads, and 3) plazas, on the other hand, are controlled by the Decentraland development team. (In particular, the Genesis plaza, where users spawn from, is in the center of the Decentraland map.) Finally, 4) districts denote a set of parcels developed by user communities based on public proposals. District founders have to obtain approval from the Decentraland DAO community to launch their district. There are currently 39 districts on the platform. [^11] https://nftplazas.com/decentraland-districts/
@@
Figure 1 shows the Decentraland map at the time of writing. We collect all the ownership transfer records, estate creation and dissolution records, and land property information until Dec. 31, 2023 from the APIs. Due to limitations on available property information, we cannot track back the addresses of districts dissolved in 2019.1111footnotemark: 11
@@ 6.1 Regression analysis
Figure 7 shows the spatial dependence of land prices ($\delta$) (see Appendix B for the details of the result). [^15] We excluded a small fraction of observations whose studentized residuals are more than 3.00 as outliers [61]. The grayed area corresponds to a 18 month window centered on 2021 Q4, which is when the hype bubble we identified earlier peaked. We first observe a clear difference between the “pre-bubble” situation from 2019 to early 2021 and the 2021–2022 bubble period. Specifically, before 2021, all prices show significant negative dependence – $\delta_{t}$ is around $-0.4$ or even lower for listing prices, indicating that sellers definitely account for location in their asking prices; buyers do, too, to a slightly lower, but nevertheless significant extent ($\delta_{t}\in[-0.2,-0.4]$). Concretely, parcels at a distance of 20 units from the Genesis Plaza ($\log D_{i}\simeq 3$) sell roughly for twice as much as the parcels at the edge of the map ($\log D_{i}\simeq 5)$.
@@ 6.2 Analysis of parcels’ value for commercial purpose
(a) The number of uploaded scenes ($N_{upload}$) normalized by the number of parcel purchases ($N_{Trade}$)
@@
Figure 9: The statistics about the building constructions on Decentraland parcels. Figure 9(a) shows the number of uploaded scenes that have social media links normalized by the number of parcel purchases. Figure 9(b) shows the number of developed parcels that have social media links normalized by the number of personally owned parcels.
@@ 7 Conclusion
This study has explored transactions on Decentraland, a decentralized virtual world platform. LAND tokens, which represent digital property title, saw their price skyrocket in 2021, which correlated with an increased amount of online discussions about Decentraland, specifically on Reddit. Topic modeling analysis revealed that discussions about Decentraland viewed that platform as representative of the NFT and metaverse ecosystems, and frequently focused on strategies to profit from these ecosystems. As a result, it is likely that Decentraland benefited from the hype surrounding NFTs and the metaverse. We performed regression and profit-and-loss analyses of land trades. We discovered that before the 2021 bubble, LAND prices followed established real-estate (bid-rent) theory, mimicking what we would see in the physical world. During the bubble, prices became disconnected from geographical considerations, leading the market to be subject to rampant short-term speculation—illustrated by the large amount of parcels being “flipped” over short periods of time. Large landowners, with multiple parcels to their names, played an important role as both buyers and sellers at the same time. In the aftermath of the bubble, new entrants were left with unrealized financial losses reaching hundreds of dollars, or they are forced to engage in a fire sale to cut off the losses.
@@
8 Acknowledgement
This research was partially supported by Ripple’s University Blockchain Research Initiative (UBRI) at Carnegie Mellon and by the Carnegie Mellon CyLab Secure Blockchain Initiative. D.K. was supported by the Japanese Government Long-Term Overseas Fellowship Program. Some of the authors hold non-negligible cryptocurrency positions, but none on the platform under study.

### v2

@@ Abstract
In the past few years, “metaverse” and “non-fungible tokens (NFT)” have become buzzwords, and the prices of related assets have exhibited large fluctuations. Are those characteristic of a speculative bubble? In this paper, we attempt to answer this question, and better understand the underlying economic dynamics. Specifically, we look at Decentraland, a virtual world platform where land parcels are sold as NFT collections. We find that initially, land prices followed traditional real estate pricing models – in particular, value decreased with distance from the most desirable areas – suggesting Decentraland behaved much like a virtual city. However, these real estate pricing models stopped applying when both the metaverse and NFTs gained increased popular attention and enthusiasm in 2021, suggesting a new driving force for the underlying asset prices. At that time, following a substantial rise in NFT market values, short-term holders of multiple parcels began to take major selling positions in the Decentraland market, which hints that, rather than building a metaverse community, early Decentraland investors preferred to cash out when land valuations became inflated. Our analysis also shows that while the majority of buyers are new entrants to the market (many of whom joined during the bubble), liquidity (i.e., parcels) was mostly provided by early adopters selling, which caused stark differences in monetary gains. Early adopters made money – more than 10,000 USD on average per parcel sold – but users who joined later typically made no profit or even incurred losses in the order of 1,000 USD per parcel. Unlike established markets such as financial and real estate markets, newly emergent digital marketplaces are mostly self-regulated. As a result, the significant financial risks we identify indicate a strong need for establishing appropriate standards of business conduct and improving user awareness.
@@ 1 Introduction
The word “metaverse” was originally coined in science fiction in 1992 [36], to denote an extension of the Internet into a single virtual world. With advances in computer graphics and virtual reality hardware, we have seen increasingly realistic attempts to build a metaverse – or at least a number of virtual worlds – over the past couple of decades, with platforms such as Second Life becoming quite popular. Tellingly, in a bid that virtual worlds were about to harness enormous economic potential, in October 2021, Facebook rebranded itself as “Meta,” hoping to become one of the main building blocks of this metaverse [45].
@@
Until recently, most virtual worlds (e.g., Second Life, World of Warcraft, etc.) were centralized, and operated by a single company. Although second-hand markets exist, in traditional, centralized virtual worlds, users willing to purchase wearables for instance to decorate their virtual-world avatars or acquire virtual land, usually have to buy them from the hosting company.
@@
On the other hand, newer projects, such as Decentraland [^1] https://decentraland.org/. or The Sandbox, [^2] https://www.sandbox.game/en/. have leveraged recent advances in decentralization technologies, e.g., “blockchains” and cryptocurrencies that instantiate bankless peer-to-peer payment systems. These newer projects rely on a blockchain to host vital records of their underlying virtual worlds. Further, governance, instead of being centralized, is handled by decentralized anonymous organizations (DAOs). Roughly speaking, DAOs attempt to implement a democratic process where governance tokens can be purchased by the general public. Similar to holding company shares, votes are typically weighted by the amount of governance tokens one holds [37]. Likewise, decentralized virtual world items (e.g., wearables, land, and so forth) are often instantiated as “non-fungible tokens” (NFTs), which are also supported by (distributed) blockchain infrastructure rather than a hosting company. Importantly, both governance tokens and NFTs can be traded outside the virtual world platform, e.g., on cryptocurrency exchanges, or third-party NFT marketplaces such as OpenSea. [^3] https://opensea.io.
@@
Decentralized virtual worlds appear to have capitalized on the surge in popularity of both the metaverse and NFTs. Concretely, in late 2021, the renewed interest in the metaverse following Facebook’s announcement, combined with the simultaneous sudden and substantial rise in cryptocurrency prices, caused a massive increase in Decentraland virtual land (parcel) prices, and in the valuation of its governance token called MANA. The median sales price of land in Decentraland rose from less than 1,000 USD in 2020 to roughly 15,000 USD in Q4 of 2021—only to plummet back to 5,000 USD in the subsequent half year. By the end of 2023, prices had gone back down to approximately 1,000 USD, and, as we will see, activity was minimal. At the time of writing, parcels go for approximately 200 USD.
@@
Such volatility shows the hallmarks of an economic bubble. Was that indeed the case? How did the situation affect the wealth of all investors participating in Decentraland—both buyers and sellers? Were certain groups of users disproportionately and/or adversely affected? If so, could regulation help “level the playing field?” Our main contribution is addressing these questions based on a quantitative analysis of online conversations on Reddit and of Ethereum blockchain records about Decentraland.
@@
Second, we examine whether traditional (physical) real estate pricing models apply to Decentraland land prices to understand if at some time these prices reflected any traditional notions of the fundamental value of real estate. Decentraland uses a finite Euclidean topology to subdivide the entire world into parcels, each of which is paired to a unique NFT. Similar to holding a property title in the physical world, the owner of that NFT owns the parcel, and, accordingly, can modify its contents (e.g., build something), start a business, or sell it. Different from the physical world, though, Decentraland allows users to instantly teleport anywhere within the virtual world. The question then becomes whether parcel prices follow traditional real-estate price patterns, with prices decreasing as a function of distance from a desirable location. This would indicate a property characteristic of a “true” virtual city.
We show that in the early period when prices were low, parcel prices do in fact appear to follow this class of traditional real-estate models quite closely. But then, during the period of the rapid rise in land prices, location no longer appears to matter much–investors seem to purchase plots like they purchase other NFTs (e.g., art or collectibles), presumably for speculative purposes. In other words, land prices begin to deviate from the traditional real-estate models, suggesting a bubble-like phenomenon.
Our analysis of market demographics further shows that two distinct groups played major roles during the period we study. New entrants (i.e., those who started trading during the bubble), most of whom appear to be individual investors, were dominantly net-buyer groups (i.e., they bought more than they sold). On the other hand, early adopters provided liquidity (i.e., parcels) to the market and were mostly selling. This caused stark differences in monetary profits: Early adopters reaped 15,000 USD worth of profit per parcel on average. On the other hand, new entrants could rarely profit from the sales – in general, they lost in the order of 1,000 USD per parcel. In short, the time at which prices were high led to a disproportionate flow of money from new participants to early adopters.
Digital item marketplaces fostered by the metaverse and NFT are currently subject to lax or ill-defined regulations [25, 40, 56]. However, our results highlight significant financial risk to new market entrants, who may have been lured by the promise of riches, when the situation was markedly unstable. These new market entrants may not have had enough information to make rational or well-informed decisions. This result echoes the negative impact of the frenzy surrounding blockchain technology [17] and emphasizes the need for establishing appropriate standards of business conduct and improving user awareness, for example, through education or mandatory risk disclosure.
@@ 2 Related work and background
Public blockchains were originally introduced to implement a secure and transparent ledger system without relying on trusted third parties [49]. However, this technology is now supporting a large volume of mostly speculative cryptocurrency trading [11, 58, 39]. Besides cryptocurrencies, blockchains are also used to support non-fungible tokens (NFTs) [48]. Smart contracts, scripts to automate transactions, make secure and transparent tracking of each token’s ownership possible, by following established NFT standards [24]. In particular, even though all transactions are executed on publicly available ledger systems, NFT ownership is resistant to tampering, without relying on trusted third-parties. Hence, NFTs have been prominently touted as a technology that facilitates the trade of creative works in a secure and transparent manner. Another desirable feature of NFTs is that, being supported by decentralized blockchains, they are platform-independent. In particular, they are not tied to specific seller marketplaces: anybody can buy and sell them. As a result, metaverse platforms that use NFTs to represent in-world items allow these items to be traded outside the platform.
Our hope for this work is to complement the literature that has evidenced the possible negative impact of hype cycles surrounding blockchain technology [17, 37], by providing a quantitative analysis of the economic dynamics of the NFT market for metaverse-related items. More broadly, we hope to contribute to the pressing agenda of better understanding emerging “Web3.0” businesses [12].
@@ 3 Dataset
This section briefly describes Decentraland first and introduces our dataset. Our measurements and analysis are based on three data sources: (1) Ethereum blockchain records, (2) Decentraland public APIs, and (3) Reddit posts.
@@ 3.1 Decentraland dataset
Decentraland is an online platform developed since 2015. Public interest in Decentraland grew in 2017 when developers held an initial coin offering (ICO) of their MANA governance token along with the first land auction. In the auction, participants had the option to bid on parcels of land or stake their MANA toward community-initiated districts. [^5] See https://medium.com/decentraland/the-terraform-event-is-around-the-corner-bdda6a2d4367 and https://github.com/decentraland/districts. Following a second land auction in 2018 and the introduction of developer tools, the official Decentraland client was launched in February 2020. At the time of writing Decentraland has published roughly 93,000 parcels of land as an NFT collection called LAND.
@@
Ether (ETH), the utility token for the Ethereum blockchain, has the second-largest market capitalization (after Bitcoin) at the time of writing. [^6] https://coinmarketcap.com/ Ethereum’s main feature is the ability to deploy smart contracts. In a nutshell, smart contracts are programs independently computed by all participants in the (Ethereum) network, who then collectively reach a consensus on the program output. This allows computations to be extremely robust to attacks. Decentraland employs smart contracts as a critical infrastructure. Because all Ethereum transactions, including all interactions with smart contracts, are recorded in a public ledger (the Ethereum blockchain), we can readily explore details of any Decentraland transaction, such as the contract functions that were called, the parties to the transaction, and any fees they might have incurred.
@@
Decentraland provides public APIs to collect parcel data, including parcel property information and ownership transfer history, sorted chronologically. [^10] https://github.com/decentraland/LAND-permissions-graph; https://github.com/decentraland/atlas-server. Every LAND token, including those bundled as EST tokens, has one or more property tags attached to it to denote the type of property the parcel refers to: 1) personally-owned lands can be bought and sold by users; 2) roads, and 3) plazas, on the other hand, are controlled by the Decentraland development team. (In particular, the Genesis plaza, where users spawn from, is in the center of the Decentraland map.) Finally, 4) districts denote a set of parcels developed by user communities based on public proposals. District founders have to obtain approval from the Decentraland DAO community to launch their district. There are currently 39 districts on the platform. A further 17 districts were dissolved in 2019.111111https://nftplazas.com/decentraland-districts/
@@
Figure 1 shows the Decentraland map at the time of writing. We collect all the ownership transfer records, estate creation and dissolution records, and land property information until Dec. 31, 2023 from the APIs. Due to limitations on available property information, we cannot track back the addresses of districts dissolved in 2019.
@@ 6.1 Regression analysis
Figure 7 shows the spatial dependence of land prices ($\delta$) (see Appendix B for details). [^15] We excluded a small fraction of observations whose studentized residuals are more than 3.00 as outliers [61]. The grayed area corresponds to a 18-month window centered on 2021 Q4, which is when the hype bubble we identified earlier peaked. We first observe a clear difference between the “pre-bubble” situation from 2019 to early 2021 and the 2021–2022 bubble period. Specifically, before 2021, all prices show significant negative dependence – $\delta_{t}$ is around $-0.4$ or even lower for listing prices, indicating that sellers definitely account for location in their asking prices; buyers do, too, to a slightly lower, but nevertheless significant extent ($\delta_{t}\in[-0.2,-0.4]$). Concretely, parcels at a distance of 20 units from the Genesis Plaza ($\log D_{i}\simeq 3$) sell roughly for twice as much as the parcels at the edge of the map ($\log D_{i}\simeq 5)$.
@@ 6.2 Analysis of parcels’ value for commercial purpose
(a) The number of uploaded scenes ($N_{upload}$) normalized by the number of parcel purchases ($N_{Trade}$).
@@
Figure 9: Building statistics on Decentraland parcels. Figure 9(a) shows the number of uploaded scenes that have social media links normalized by the number of parcel purchases. Figure 9(b) shows the number of developed parcels that have social media links normalized by the number of personally owned parcels.
@@ 7 Conclusion
This study has explored transactions on Decentraland, a decentralized virtual world platform. LAND tokens, which represent digital property title, saw their price skyrocket in 2021, which correlated with an increased amount of online discussions about Decentraland, specifically on Reddit. Topic modeling analysis revealed that discussions about Decentraland viewed that platform as representative of the NFT and metaverse ecosystems, and frequently focused on strategies to profit from these ecosystems. As a result, it is likely that Decentraland benefited from the hype surrounding NFTs and the metaverse. We performed regression and profit-and-loss analyses of land trades. We discovered that before the 2021 hype cycle, LAND prices followed established real-estate (bid-rent) theory, mimicking what we would see in the physical world. During that hype cycle, prices became disconnected from geographical considerations, leading the market to be subject to rampant short-term speculation—illustrated by the large amount of parcels being “flipped” over short periods of time, and characteristic of an economic bubble. Large landowners, with multiple parcels to their names, played an important role as both buyers and sellers at the same time. In the aftermath of the bubble, new entrants were left with unrealized financial losses reaching hundreds of dollars, or were forced to engage in fire sales to cut off their losses.
@@
8 Acknowledgements
This research was partially supported by Ripple’s University Blockchain Research Initiative (UBRI) at Carnegie Mellon and by the Carnegie Mellon CyLab Secure Blockchain Initiative. D.K. was supported by the Japanese Government Long-Term Overseas Fellowship Program. Some of the authors hold non-negligible cryptocurrency positions, but none on Decentraland. As part of exploratory efforts, some of the authors acquired a negligible amount of Decentraland real estate.